OpenStudy (anonymous):

use rational exponents to simplify 4sqrtx^20 My answer is x^5 is that right?

OpenStudy (anonymous):

Im trying to put the equation like it is on my paper but I can't

OpenStudy (anonymous):

\[\sqrt[4]{x ^{20}}\]

OpenStudy (anonymous):

\[=x ^{20/4}\]

OpenStudy (anonymous):

\[=x ^{5}\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the problem 4\[\sqrt{x^{20}}\]

OpenStudy (anonymous):

with the 4 on the outside

OpenStudy (anonymous):

on the left side, sry

OpenStudy (anonymous):

so yes it will be like this \[\sqrt[4]{x ^{20}}\]

OpenStudy (anonymous):

yes like that, so will the answer still be x^5

OpenStudy (anonymous):

yes \[\sqrt[4]{x ^{20}} = x^{20/4} = x^{5}\]

OpenStudy (anonymous):

oh ok, thank you :)